OpenStudy (anonymous):

solve by factoring 49x^2 - 16 = 0

OpenStudy (paystub-2):

Is this your very first ques. on here?

OpenStudy (anonymous):

this = (7x-4)(7x+4) do you need anything beyond that?

OpenStudy (anonymous):

yes :/

OpenStudy (anonymous):

i think that should help. Thank you :)

OpenStudy (anonymous):

x= 4/7, -4/7

OpenStudy (anonymous):

This is what's called a perfect square, which means there isn't a middle term. \[49x ^{2}-16=0\] (7x-4)(7x+4)=0 Now check using FOIL: First: 7x X 7x= \[49x^{2}\] Outside: 7x X 4= 28x Inside: 7x X -4= -28x Last: -4 X 4= -16 The 28x and -28x cancel each other out and you are left with \[49x ^{2}-16\]